OpenStudy (ohhskyzzx3):

Really Need Help. Will fan and give medal!! Let f(x)=sqrt of x-2 and g(x) = sqrt of x+7. Find (f*g)(x). Assume all appropiate restrictions to the domain.

OpenStudy (ohhskyzzx3):

Answer Choices: (f*g)(x) = x^2 + 5x - 14 (f*g)(x) = x^2 + 9x - 14 (f*g)(x) = sqrt of (x^2 + 5x - 14) (f*g)(x) = sqt of (x^2 + 9x - 14)

OpenStudy (dunk789):

B

OpenStudy (ohhskyzzx3):

Are you sure?

OpenStudy (dunk789):

no

OpenStudy (dunk789):

try using mathway

OpenStudy (ohhskyzzx3):

ok

OpenStudy (ohhskyzzx3):

The answer is C everybody

OpenStudy (retireed):

I finally proved it to myself by graphing the two functions. It is only true for the domain x>0